OpenStudy (anonymous):

Suppose a plane drops water from a height of 1,024 feet. How many seconds will it take for the water to hit the ground? 4 16 9 8

OpenStudy (anonymous):

4

OpenStudy (anonymous):

How do you figure that out?

OpenStudy (anonymous):

@KayleesaKay

OpenStudy (anonymous):

I'm not sure, I just used logical reasoning. Sorry :(

OpenStudy (anonymous):

@Mertsj Help please?

OpenStudy (mertsj):

Don't you have a formula or equation given?

OpenStudy (anonymous):

Oh yeah, I have the question before it I guess; Airplanes are often used to drop water on forest fires in an effort to stop the spread of the fire. The time t it takes the water to travel from height h to the ground can be derived from the equation h = 1/2 • gt2 where g is the acceleration due to gravity (32 feet/second2).

OpenStudy (mertsj):

h=1/2gt^2 1024=1/2(32)t^2

OpenStudy (mertsj):

Can you solve that?

OpenStudy (anonymous):

I got 8, is that right?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw